OpenStudy (anonymous):

A malted milk ball with a radius of ¼" (0.25 in) is coated with chocolate. What is the surface area of the chocolate? Ignore the thickness. Leave your answer in terms of π.

zepdrix (zepdrix):

The surface area of a sphere is given by:\[\Large A=4\pi r^2\] In this problem our radius is given to be \(\large r=\dfrac{1}{4}\). So just plug that in and solve for A! :) They told us they want the answer in `terms of pi`. So do not multiply out the pi as 3.14 or anything like that. Leave it as \(\large \pi\). \[\Large A=4\pi \left(\frac{1}{4}\right)^2\]

OpenStudy (anonymous):

I got .7853

zepdrix (zepdrix):

So when we square 1/4 we get,\[\Large A=4\pi\left(\frac{1}{16}\right)\]Let's multiply the 4 and 1/16,\[\Large A=\frac{4}{16}\pi \qquad\to\qquad A=\frac{1}{4}\pi\]This is in terms of pi. See how I left the \(\large \pi\) icon? You got the correct answer, but don't multiply out the pi to give you a decimal! :)

OpenStudy (anonymous):

So it would just be \[A=1/4\Pi \] ?

zepdrix (zepdrix):

yes :)

OpenStudy (anonymous):

Awesome! Thanks! You rock!!!(: